# Gaussian On-the-Fly Splatting: A Progressive Framework for Robust Near Real-Time 3DGS Optimization

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) achieves high-fidelity rendering with fast real-time performance, but existing methods rely on offline training after full Structure-from-Motion (SfM) processing. In contrast, this work introduces On-the-Fly GS, a progressive framework enabling near real-time 3DGS optimization during image capture. As each image arrives, its pose and sparse points are updated via on-the-fly SfM, and newly optimized Gaussians are immediately integrated into the 3DGS field. We propose a progressive local optimization strategy to prioritize new images and their neighbors by their corresponding overlapping relationship, allowing the new image and its overlapping images to get more training. To further stabilize training across old and new images, an adaptive learning rate schedule balances the iterations and the learning rate. Moreover, to maintain overall quality of the 3DGS field, an efficient global optimization scheme prevents overfitting to the newly added images. Experiments on multiple benchmark datasets show that our On-the-Fly GS reduces training time significantly, optimizing each new image in seconds with minimal rendering loss, offering the first practical step toward rapid, progressive 3DGS reconstruction. https://xywjohn.github.io/GS_On-the-Fly.github.io/
@@ I INTRODUCTION
Novel View Synthesis (NVS) aims to generate realistic images of novel views based on a set of input images[1]. Recently, Kerbl et al.[3] introduced a novel 3D representation technique known as 3D Gaussian Splatting (3DGS), which employs a large number of 3D Gaussian kernels to represent a 3D scene. Subsequent research has made notable advancements in improving the 3DGS field in terms of training efficiency[4, 5, 6, 7, 8, 9, 10, 11, 12], rendering speed[13, 14], and output quality[15, 16, 17, 18, 19]. However, most existing approaches rely on an offline training solution, which means that 3DGS can only be optimized after all images are processed by Colmap[35], preventing immediate synthesis of novel views upon completion of data acquisition[20]. While some recent works [20, 21, 39, 40, 41] have made progress towards real-time/near real-time training of 3DGS by integrating SLAM, such as GS-SLAM[21], MonoGS++[41]. They take indoor video frames with spatial and temporal continuity as input, which might degenerate in some practical applications[36, 38].
To address these limitations, we propose On-the-Fly GS, a progressive framework for near real-time 3DGS field training that enables 3D Gaussian training during discrete image capturing. As shown in Fig. 1, in our work, a new image can be acquired in an arbitrary manner. Its pose and relevant new point cloud are estimated online by the on-the-fly SfM[38] [^1] See more at https://yifeiyu225.github.io/on-the-flySfMv2.github.io/, which are then inputted into our on-the-fly GS to optimize the corresponding local 3DGS field in near real time and update the entire 3DGS field.
@@
In contrast to the original 3DGS that focuses on reconstructing a static scenario using a fixed set of images and their corresponding poses, our On-the-Fly GS targets a dynamically expanding scenario, where images are progressively captured over time. This shift necessitates a different training strategy tailored for the incremental expansion of the 3DGS fields. A key challenge in training dynamic scenes lies in the inconsistency of local quality within the 3DGS fields. Images captured earlier undergo more training iterations, resulting in higher rendering quality, while newly acquired images, having undergone fewer training iterations, exhibit lower rendering quality. Consequently, the learning rate decay strategy of the original 3DGS based on complete 3DGS field quality is no longer applicable. Instead, On-the-Fly GS calculates the learning rate for each image based on the number of iterations it has undergone. Images with fewer training iterations are assigned higher learning rates, while images with more iterations receive lower learning rates.
Furthermore, in the local optimization, based on the overlapping relationships, each local connected image is assigned a weight that determines the number of training iterations it will undergo. This weight encourages prioritizing training the newly added images and their local overlapping images, thereby accelerating the rendering quality in regions captured more recently. To further enhance training efficiency and near real-time performance for each image, two strategies have been integrated into On-the-Fly GS: inspired by Adr-Gaussian[5], a load-balancing strategy is incorporated to mitigate the over-accumulation of 3D Gaussians, while simultaneously ensuring fast training of newly added regions; Drawing from insights from Taming-3DGS[9], we re-implement the back-propagation by transitioning from pixel-based parallelism to splatting-based parallel computation. In summary, this work makes the following contributions:
@@
  • To achieve robust and near real-time progressive 3DGS training, self-adaptive methods are presented to determine the training iteration number of each image and to update the learning rate during each iteration.
  • Our On-the-Fly GS reduces training time significantly on several benchmark datasets, offering the first practical step toward rapid 3DGS optimization during image capturing.
@@ II RELATED WORKS
Novel View Synthesis (NVS). Novel View Synthesis has been a central research topic in both computer vision and computer graphics for many years[22]. Traditionally, various approaches have been developed to achieve NVS. Light Field Rendering employs light field cameras to capture light field data, enabling novel view synthesis by interpolation[23, 24]. The Lumigraph method [25, 26] improves Light Field Rendering with greater flexibility and broader applicability. View Interpolation [27] utilizes multi-view images to interpolate among existing images, while it still employs a traditional photogrammetry method to generate a mesh model of the scene and renders images from arbitrary viewpoints using back-projection[2]. In recent years, the advent of Neural Radiance Fields (NeRF)[28] and 3D Gaussian Splatting (3DGS)[3] has introduced novel paradigms for NVS. NeRF, based on volumetric rendering, leverages a multi-layer perceptron (MLP) to implicitly represent a 3D scene. The original NeRF demonstrated remarkable performance in small-scale scenes (e.g., desk-size objects), subsequent NeRF variants further improved its performance by reducing training time[29], enhancing rendering efficiency[30] and quality[31], enabling large-scale scene modeling[32], etc. However, contemporary NeRF methods still face significant challenges in complex, unbounded outdoor scenes where achieving high-quality, high-resolution rendering often requires extensive training time, making real-time rendering infeasible. To address these limitations, Kerbl et al.[3] introduced 3DGS to explicitly represent 3D scenes.
@@
3DGS fast training. Although 3DGS delivers high-quality and real-time rendering, its training process remains computationally intensive. Consequently, extensive research has been dedicated to accelerating the 3D Gaussians optimization. Several studies have sought to expedite 3DGS training by reducing the number of Gaussians in the scene representation[4, 5, 6, 7, 8]. Panagiotis et al.[4] proposed an efficient and resolution-aware Gaussians pruning strategy along with an adaptive adjustment method for the degree of spherical harmonic coefficients. It reduces the number of Gaussians by more than 50% and saves the memory. Adr-Gaussian[5] introduced a load balancing strategy to optimize the distribution of Gaussians and prevent their excessive accumulation in specific regions, thereby reducing the number of Gaussians. EfficientGS[7] selectively restricted the increase in key Gaussians and implemented a pruning mechanism to remove redundant ones, retaining critical Gaussians while removing unnecessary ones. Similarly, LightGaussian[8] introduced a global importance-based pruning strategy, retaining critical and removing redundant Gaussians. Since 3DGS training requires optimizing the parameters of a large number of Gaussians, reducing 3D Gaussians can significantly enhance training speed and decrease memory overhead.
@@
Some studies also focused on refining training strategies to improve computational efficiency. Adr-Gaussian[5] adopts an early culling strategy during forward-propagation, refining the influence range of each Gaussian. Taming 3DGS[9] implements a Gaussian-based parallel strategy for gradient accumulation and optimizes the computation of spherical harmonics and differentiable loss functions, thereby enhancing back-propagation efficiency. RetinaGS[10] and DOGS[11] address the challenge of large-scale scenes and high-resolution image data. They propose a parallel training method that divides the whole scene into several overlapping blocks, enabling parallel and accelerating training for large-scale scenes. Unlike methods that focus on reducing the number of Gaussians, these approaches optimize specific computational steps in the training pipeline, achieving significant speed improvements.
3DGS real-time training. Recent studies have explored real-time training of 3DGS. Before 3DGS, through integration with SLAM, NICE-SLAM[39] and Vox-Fusion[40] successfully achieve real-time NeRF training. In addition, Gaussian Splatting SLAM[21], MonoGS++[41] and RTG-SLAM[20] have further improved real-time 3DGS training, delivering higher rendering quality. While these methods enable real-time 3DGS training, they struggle to handle discrete images that are not with spatially and temporally continuous. In contrast, our On-the-Fly GS, which can deal with images captured in an arbitrary way, employs a progressive framework for near real-time 3DGS training. It addresses the limitations of existing SLAM-based methods and extends the applicability of 3DGS training to more general use cases.
@@ III PRELIMINARIES
3D Gaussian Splatting. 3DGS employs a large number of Gaussians to represent a real 3D scene[3]. The geometric properties of each Gaussian are defined using a 3D Gaussian distribution:
$G(X)=e^{-\frac{1}{2}\left(X-X_{0}\right)^{T}\boldsymbol{\Sigma}^{-1}\left(X-X_{0}\right)},$ (1)1( 1 )
Where $X_{0}$ represents the center position of a Gaussian, $X$ represents the position of a sampled point, and $\Sigma$ corresponds to the covariance matrix of the Gaussian distribution. For each Gaussian, its covariance matrix $\Sigma$ can be formulated by a rotation matrix $R$ and a scaling matrix $S$, as follows:
$\Sigma=RSS^{T}R^{T}.$ (2)2( 2 )
The above parameters define the geometric properties of a Gaussian. Besides, 3DGS uses spherical harmonics (SH) and opacity $\sigma$ to reveal the material properties. Before calculating the color of one specific pixel, 3DGS estimates the splatting opacity $\alpha$ based on the Gaussian-pixel 2D distance $x$ and the projected covariance matrix $\Sigma^{\prime}$:
$\alpha=\sigma e^{-\frac{1}{2}x^{T}\Sigma^{{}^{\prime}-1}x}.$ (3)3( 3 )
Given the splatting opacities $\alpha_{i}$ and color $c_{i}$ calculated from SH of $N$ Gaussians associated with the current pixel, 3DGS calculates the color of the current pixel $c_{p}$ by alpha blending, as follows:
$c_{\text{pixel}}=\sum_{i=1}^{N}\left(c_{i}a_{i}\prod_{j=1}^{i-1}\left(1-a_{j}\right)\right)$ (4).4( 4 ) .
During a novel view rendering, 3DGS assigns a dedicated thread to each pixel, enabling parallel computation of the color of each pixel to facilitate high-quality, real-time image rendering.
On-the-Fly SfM. Before training the original 3DGS field, COLMAP[35] is typically used for pose estimation and sparse point cloud generation. However, COLMAP requires feature matching across all images, making it impractical in scenarios where image acquisition and 3DGS training occur simultaneously, such as our On-the-Fly GS training. To overcome this limitation, instead of COLMAP, we use the On-the-Fly SfM[38] for our subsequent experiments. On-the-Fly SfM employs a Hierarchical Weighted Local Bundle Adjustment (BA), in which only neighboring connected and registered images are incorporated into BA after acquiring a new image. Unlike COLMAP’s BA, which incurs a substantial computational cost as the number of images increases, Hierarchical Weighted Local BA significantly reduces time complexity, ensuring efficient processing. Additionally, On-the-Fly SfM can achieve online feature matching via global feature retrieval on HNSW (Hierarchical Navigable Small World)[43], enabling efficient dynamic updates for the matching matrix. On-the-fly SfM allows for near real-time pose estimation and sparse point cloud generation immediately after capturing a new image, ensuring a continuous and timely input for our approach.
@@ IV-A Overview of on-the-fly GS
To achieve near real-time 3DGS field optimization, the 3DGS field is supposed to be trained concurrently with image acquisition. Therefore, rather than relying on the original offline training solution of 3DGS, we propose a progressive training framework. As illustrated in Fig. 2, our training process is structured into three distinct phases:
Figure 2: Workflow of the proposed On-the-Fly GS.
Phase 1: Initial 3DGS Training. Training the 3DGS field with a limited number of images risks overfitting to the initial few images (e.g., 2 or 3 images)[34]. To mitigate this, we begin by training an initial 3DGS field after On-the-Fly SfM has estimated the poses and generated the sparse point cloud for $N_{0}$ initial images. This initial 3DGS field serves as the foundation for subsequent progressive training. In this phase, we adopt the original 3DGS training method but set a more severe number for training iterations to prevent overfitting to these initial images.
Phase 2: Progressive 3DGS Training. As a new image is acquired, On-the-Fly SfM continuously updates the camera poses of both the newly acquired and previously registered images, while refining the sparse point cloud. Based on the updated sparse point cloud, some new Gaussians are appended to the previous 3DGS field. The updated model will be then optimized. The optimization process contains local optimization and global optimization. The local optimization focuses on fast optimization of the newly added regions of the 3DGS field, while the global optimization ensures the overall quality and consistency of the 3DGS field and prevents overfitting to the newly added images.
Phase 3: Final Optimization. Once all images have been acquired and the progressive training is complete, a final optimization is performed on the current 3DGS field. This step further refines the model and enhances its overall quality, ensuring the best reconstruction result.
@@ IV-C Self-adaptive Training Iteration Allocation via Hierarchical Image Weighting
Newly added images often introduce novel unexplored regions or correspond to areas where the current 3DGS field exhibits poor rendering quality. To ensure that more training is allocated to the new image and its neighboring connected images during local optimization, we propose an image hierarchical weighting strategy, which can be used to adaptively allocate training iterations for each image. The weight is determined by the overlap relationship between the newly fly-in image and the already registered images, as depicted in Fig. 3.
Figure 3: The Hierarchical Image Weighting.
Assuming that there are $n-1$ registered images, when the image $n$ ($I_{11}$, $l=1$) is newly added, the weight of the newly added image is fixed as $\omega_{11}=1$. The weights of the remaining registered images are calculated based on the match matrix online estimated by On-the-Fly SfM. If the overlapping degree of any two images is $M_{i,j}$, for images $I_{2i}$ that have a direct overlapping relationship ($M_{i,n}\neq 0,l=2$) with the newly added image, their weights $\omega_{2i}$ are calculated using the following equation:
$\omega_{2i}=\omega_{11}\times M_{i,n}=M_{i,n}.$ (5)5( 5 )
For images $I_{li}$ that have an indirect overlapping relationship ($M_{i,n}=0,l\geq 3$) with the newly added image, their weights are determined by their overlapping relationship with images in the previous layer and the weights of those images. If the layer $k-1$ contains $D$ images, the weight of the image $j$ in the layer $k-1$ is $\omega_{(k-1)j}$, the weight of the corresponding image $i$ in layer $k$ will be calculated as follows:
$\omega_{ki}=\frac{{\textstyle\sum_{j=1}^{D}}\omega_{\left(k-1\right)j}\times M_{\left(k-1\right)j,ki}}{D}$ (6)6( 6 )
Based on the hierarchical weighting method, we can adaptively assign a weight to all registered images including the newly added image. Subsequently, the training iterations for local neighboring images will be allocated based on their weights. For $N$ images, with given weight $\omega_{i}$ of image $i$, the training iterations $T_{i}$ can be calculated as follows:
$T_{i}=\frac{\omega_{i}}{{\textstyle\sum_{j=1}^{N}}\omega_{j}}\times\text{LOT}$ (7)7( 7 )
LOT represents the number of training iterations required during each local optimization. This hierarchical weighting method assigns higher weights for the newly added image and its neighboring connected images while the remaining images are given lower weights, ensuring the newly added image and its neighboring connected images receive more training in the current local optimization.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) achieves high-fidelity rendering with fast real-time performance, but existing methods rely on offline training after full Structure-from-Motion (SfM) processing. In contrast, this work introduces Gaussian on-the-fly Splatting (abbreviated as On-the-Fly GS), a progressive framework enabling near real-time 3DGS optimization during image capture. As each image arrives, its pose and sparse points are updated via on-the-fly SfM[1] [^1] See more at https://yifeiyu225.github.io/on-the-flySfMv2.github.io/, and newly optimized Gaussians are immediately integrated into the 3DGS field. To achieve this, we propose a progressive Local & Semi-Global optimization to prioritize the new image and its neighbors by their corresponding overlapping relationship, allowing the new image and its overlapping images to get more training. To further stabilize training across previous and new images, an adaptive learning rate schedule balances the iterations and the learning rate. Extensive experiments on multiple benchmarks show that our On-the-Fly GS reduces training time significantly, optimizing each new image in seconds with minimal rendering loss, offering one of the first practical steps toward rapid, progressive 3DGS reconstruction.
@@ I INTRODUCTION
Novel View Synthesis (NVS) aims to generate realistic images of novel views based on a set of input images[2]. Recently, Kerbl et al.[3] introduced a novel 3D representation technique known as 3D Gaussian Splatting (3DGS), which employs a large number of 3D Gaussian kernels to represent a 3D scene. Subsequent research has made notable advancements in improving the 3DGS field in terms of training efficiency[4, 5, 6, 7, 8, 9], rendering speed[10, 11], and output quality[12, 13, 14, 15, 16]. However, most existing approaches rely on an offline training solution, which means that 3DGS can only be optimized after all images are processed by COLMAP[17], preventing immediate synthesis of novel views upon completion of data acquisition[18]. Some recent works[18, 19, 20, 21] have made progress towards real-time/near real-time training of 3DGS by integrating SLAM, such as GS-SLAM[19], RTG-SLAM[18]. They take indoor video frames with inherent spatial and temporal continuity as input, which might degenerate in some practical applications[22, 1]. More detailed differences between SLAM-based 3DGS methods and our On-the-Fly GS are discussed in Sect.II.
To address these limitations, we propose On-the-Fly GS, a progressive framework for near real-time 3DGS field training that enables 3D Gaussian training during discrete image capturing. As shown in Fig. 1, in our work, a new image can be acquired in an arbitrary manner. Its pose and relevant new point clouds are estimated online by the On-the-Fly SfM[1], which are then input into our On-the-Fly GS to optimize the corresponding local 3DGS field in near real time and update the entire 3DGS field.
@@
In contrast to the original 3DGS, which focuses on reconstructing a static scenario using a fixed set of images and their corresponding poses, our On-the-Fly GS targets a dynamically expanding scenario, where images are progressively captured over time. This shift necessitates a different training strategy tailored for the incremental expansion of the 3DGS field. One key challenge in training dynamically extended scenes lies in the inconsistency of local quality within the 3DGS field. Images captured earlier undergo more training iterations, resulting in higher rendering quality, while newly acquired images, having undergone fewer training iterations, exhibit lower rendering quality. Consequently, the learning rate decay strategy of the original 3DGS based on complete 3DGS field quality is no longer applicable. Instead, On-the-Fly GS calculates the learning rate for each image based on the number of iterations it has undergone. Images with fewer training iterations are assigned higher learning rates, while images with more iterations receive lower learning rates.
Furthermore, in the Local & Semi-Global optimization, based on the overlapping relationships, each local connected image is assigned a weight that determines the number of training iterations it will undergo. This weight encourages prioritizing training the newly added images and their local overlapping images, thereby accelerating the rendering quality in regions captured more recently. To further enhance training efficiency and near real-time performance for each image, two strategies have been integrated into On-the-Fly GS: inspired by Adr-Gaussian[5], a load-balancing strategy is incorporated to mitigate the over-accumulation of 3D Gaussians, while simultaneously ensuring fast training of newly added regions; Drawing from insights from Taming-3DGS[7], we re-implement the back-propagation by transitioning from pixel-based parallelism to splatting-based parallel computation. In summary, this work makes the following contributions:
@@
  • To achieve robust and near real-time progressive 3DGS training, a Local & Semi-Global optimization strategy is designed, and self-adaptive methods are presented to determine the training iteration number of each image and to update the learning rate during each iteration.
  • Our On-the-Fly GS reduces training time significantly on several benchmark datasets, offering one of the first practical steps toward rapid 3DGS optimization during image capturing.
@@ II RELATED WORKS
Novel View Synthesis (NVS). Traditional photogrammetry method takes unposed images as input data, generates a mesh model of the current scene through a complex pipeline, and renders images from arbitrary viewpoints using back-projection[23]. In recent years, the advent of Neural Radiance Fields (NeRF)[24] and 3D Gaussian Splatting (3DGS)[3] has introduced novel paradigms for NVS. NeRF, based on volumetric rendering, leverages a multi-layer perceptron (MLP) to implicitly represent a 3D scene. Although NeRF has demonstrated strong rendering capabilities, it still requires substantial computational resources when facing significant challenges in complex, unbounded outdoor scenes, where achieving high-quality, high-resolution rendering. To address these limitations, Kerbl et al.[3] introduced 3DGS to explicitly represent 3D scenes. 3DGS offers faster training and rendering speeds for novel views while maintaining acceptable rendering quality, which has led to widespread attention since its introduction.
3DGS fast training. Although 3DGS delivers high-quality and real-time rendering, its training process remains computationally intensive, making it difficult to directly meet the demands of near real-time or real-time applications. Consequently, extensive research has been dedicated to accelerating the 3DGS optimization.
@@
Several studies have sought to expedite 3DGS training by reducing the number of Gaussians in the scene representation. Panagiotis et al.[4] proposed an efficient and resolution-aware Gaussian pruning strategy along with an adaptive adjustment method for the degree of spherical harmonic coefficients. It reduces the number of Gaussians by more than 50% and saves memory. EfficientGS[6] and LightGaussian[25] evaluate the contribution of each Gaussian, only retaining the critical Gaussians and removing the redundant ones. Adr-Gaussian[5] introduced a load balancing strategy to optimize the distribution of Gaussians and prevent their excessive accumulation in specific regions while also avoiding too few Gaussians in other areas. Since 3DGS training requires optimizing the parameters of a large number of Gaussians, reducing 3D Gaussians can significantly enhance training speed and decrease memory overhead.
@@
Unlike methods that focus on reducing the number of Gaussians, some studies also focused on refining training strategies to improve computational efficiency. Taming 3DGS[7] implements a Gaussian-based parallel strategy for gradient accumulation and optimizes the computation of spherical harmonics and differentiable loss functions, thereby enhancing back-propagation efficiency. Facing the challenge of large-scale scenes and high-resolution image data, RetinaGS[8] and DOGS[9] propose a parallel training method that divides the whole scene into several overlapping blocks, enabling parallel and accelerating training for large-scale scenes.
3DGS real-time training. Recent studies have explored real-time training of 3DGS. Before 3DGS, through integration with SLAM, NICE-SLAMM[20] and Vox-Fusion[21] successfully achieved real-time NeRF training. More recently, 3DGS has also achieved real-time training via similar integration with SLAM. GS-SLAM[19] predicts and continuously refines the poses of keyframes. Based on depth priors, it adds a certain number of Gaussians to the 3DGS field, only focusing the optimization primarily on newly added and recently added keyframes. RTG-SLAM[18] also leverages depth to support 3DGS training and further improves the strategies for both Gaussian addition and optimization. Compared to GS-SLAM, it achieves significantly better training efficiency and novel view rendering quality. Subsequently, numerous efforts have been made to further improve SLAM-based 3DGS methods. Deng et al.[26] propose a sliding window-based masking strategy to reduce redundant Gaussians, thereby improving training efficiency and reducing memory consumption. GSFusion[27] incorporates each Gaussian into a volumetric mapping system to leverage geometric information, significantly accelerating the training process while maintaining high rendering quality. However, their reliance on depth sensors would be limited when dealing with large-scale outdoor scenes or UAV images. Although some existing methods take RGB-only data as input (such as GS-SLAM[19], WildGS-SLAM[28], DROID-Splat[29], On-the-Fly NVS[30]), they still struggle to handle discrete images that are not spatially and temporally continuous.
In contrast, our On-the-Fly GS, which can deal with images captured in an arbitrary way without spatiotemporal continuity, employs a progressive framework for near real-time 3DGS training. It only requires that the newly fly-in image has a certain degree of overlap with all previously registered images, and does not rely on depth priors for 3DGS field expansion. It addresses the limitations of existing SLAM-based 3DGS methods and extends the applicability of 3DGS training to more general use cases.
@@ III PRELIMINARIES
Figure 2: Workflow of the proposed On-the-Fly GS. We first initialize the 3DGS field. Then, for each new fly-in image, we update the current 3DGS field and perform the optimization on it. Finally, after a final global BA of previous SfM results, we further refine the 3DGS field to enhance rendering quality.
3D Gaussian Splatting. 3DGS employs a large number of Gaussians to represent a real 3D scene[3]. Given the rotation matrix $R$ and the scaling matrix $S$, the geometric properties of each Gaussian are defined by a covariance matrix $\Sigma=RSS^{T}R^{T}$ centered at point $X_{0}$:
$G(X)=e^{-\frac{1}{2}\left(X-X_{0}\right)^{T}\boldsymbol{\Sigma}^{-1}\left(X-X_{0}\right)}$ (1)( 1 )
Furthermore, 3DGS uses spherical harmonics (SH) and opacity $\sigma$ to reveal the material properties. Given the splatting opacities $\alpha_{i}$ estimated based on Gaussian-to-Pixel distance and color $c_{i}$ calculated from SH of $N$ Gaussians associated with the specific pixel, 3DGS calculates the color of the pixel $c_{p}$ by alpha blending, as follows:
$c_{\text{pixel}}=\sum_{i=1}^{N}\left(c_{i}a_{i}\prod_{j=1}^{i-1}\left(1-a_{j}\right)\right)$ (2)( 2 )
On-the-Fly SfM. Traditional SfM frameworks (e.g., COLMAP[17]) are incapable of processing images in an online manner. To overcome this limitation, we rely on the On-the-Fly SfM[1] for our subsequent implementation. On-the-Fly SfM employs a Hierarchical Weighted Local Bundle Adjustment (BA), in which only neighboring connected and registered images are incorporated into BA after acquiring a new image. Unlike COLMAP’s BA, which incurs a substantial computational cost as the number of images increases, Hierarchical Weighted Local BA significantly reduces time complexity, ensuring efficient processing. Additionally, On-the-Fly SfM can achieve online feature matching via global feature retrieval on HNSW (Hierarchical Navigable Small World)[31], enabling efficient dynamic updates for the matching matrix. On-the-fly SfM allows for near real-time pose estimation and sparse point cloud generation immediately after capturing a new image, ensuring a continuous and timely input for our approach.
@@ IV-A Overview of on-the-fly GS
To achieve near real-time 3DGS field optimization, we propose a progressive training framework. As illustrated in Fig. 2, our training process is structured into three distinct phases:
Phase 1: Initial 3DGS Training. Training the 3DGS field with a limited number of images risks overfitting to the initial few images (e.g., 2 or 3 images)[32]. To mitigate this, we begin by training an initial 3DGS field after On-the-Fly SfM has estimated the poses and generated the sparse point cloud for $N_{0}$ initial images. This initial 3DGS field serves as the foundation for subsequent progressive training. In this phase, we adopt the original 3DGS training method, but a stricter limit was imposed on the number of training iterations to prevent overfitting to these initial images.
Phase 2: Progressive 3DGS Training. As a new image is acquired, On-the-Fly SfM continuously updates the camera poses of both the newly acquired and previously registered images, while refining the sparse point cloud. Based on the updated sparse point cloud, some new Gaussians are appended to the previous 3DGS field. The updated 3DGS field will then be optimized. During the optimization, the new fly-in image and its neighboring images are prioritized for training to accelerate the optimization of newly expanded regions. Meanwhile, some of the other images are randomly selected as training targets to avoid degrading the entire scene, including the very previous 3DGS field.
Phase 3: Final Optimization. Once all images have been acquired and the progressive training is complete, On-the-Fly SfM further performs global BA on the estimated poses. Subsequently, a rapid additional optimization is applied to the current 3DGS field, enabling a rapid improvement in overall rendering quality within a short period of time.
@@ IV-C Self-adaptive Training Iteration Allocation via Hierarchical Image Weighting
Hierarchical Image Weighting. Newly added images often introduce novel, unexplored regions or correspond to areas where the current 3DGS field exhibits poor rendering quality. To ensure that more training is allocated to the new image and its neighboring connected images during optimization, we propose an image hierarchical weighting strategy, which can be used to adaptively allocate training iterations for each image. The weight is determined by the overlap relationship between the newly fly-in image and the already registered images, as depicted in Fig. 3.
Figure 3: Hierarchical Image Weighting. The already solved images are clustered into hierarchical layers using their overlap relationships, assigning higher weights to those that exhibit closer relationship with the new image.
Assuming that there are $n-1$ registered images, when the image $n$ ($I_{11}$, $l=1$) is newly added, the weight of the newly added image is fixed as $\omega_{11}=1$. The weights of the remaining registered images are calculated based on the match matrix online, estimated by On-the-Fly SfM. In the match matrix, the overlapping degree of any two images is calculated based on the number of matched feature points shared between them. If the overlapping degree of any two images is $M_{i,j}$, for images $I_{2i}$ that have a direct overlapping relationship ($M_{i,n}\neq 0,l=2$) with the newly fly-in image, their weights $\omega_{2i}$ are calculated using the following equation:
$\omega_{2i}=\omega_{11}\times M_{i,n}=M_{i,n}.$ (3)( 3 )
For images $I_{li}$ that have an indirect overlapping relationship ($M_{i,n}=0,l\geq 3$) with the newly added image, their weights are determined by their overlapping relationship with images in the previous layer and the weights of those images. If the layer $k-1$ contains $N_{k-1}$ images, the weight of the image $j$ in the layer $k-1$ is $\omega_{(k-1)j}$, the weight of the corresponding image $i$ in layer $k$ will be calculated as follows:
$\omega_{ki}=\frac{{\textstyle\sum_{j=1}^{N_{k-1}}}\omega_{\left(k-1\right)j}\times M_{\left(k-1\right)j,ki}}{N_{k-1}}$ (4)( 4 )
Local & Semi-Global optimization. After the addition of any new fly-in image, we perform $T_{I}$ training iterations to optimize the current 3DGS field. The training process is structured into two parts of local and semi-global ones, with each performing $T_{I}$/2 training iterations. 1) Local optimization: The top $N_{m}$ images with the highest weights will be selected as key training images. For one key training image $i$ with corresponding weight $\omega_{k}(k=1,2,...,N_{m}$), it is supposed to be trained for $T_{i}^{m}$ iterations, which is determined by Eq 5. 2) Semi-Global optimization: For the other images with low weights, the remaining half training iterations are evenly distributed among these images. This strategy ensures that more training is performed on the new fly-in image and its neighbor images, while avoiding forgetting the 3DGS field of the very previous scene.
$T_{i}^{m}=\frac{1-e^{-\omega_{i}}}{2\sum_{j=1}^{N_{m}}\left(1-e^{-\omega_{j}}\right)}\times T_{I}$ (5)( 5 )